# High Dimensional Ensemble Kalman Filter

### v1

@@ 4.1 Consistent estimation of high dimensional forecast error covariance matrix
The covariance matrix class suitable for the banding estimator is the following bandable covariance class (Bickel and Levina, 2008a)
@@
$\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}$ $\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0$
@@
for some positive $\alpha$, C and $\varepsilon_{0}$. The bandable class prescribes a decaying pattern of $|\sigma_{ij}|$ as $|i-j|$ increases with the decaying rate $\alpha$. This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in atmospheric and geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger. Note that $\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|$ is the $L_{1,1}$ norm of $B_{k}\left(\boldsymbol{\Sigma}\right)-\boldsymbol{\Sigma}$, which bounds the spectral norm from above (Golub and Van Loan, 2013).
In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or the poles of the earth, the state vectors are highly dependent for the variable components between the two ends of the state vector. This leads to a circular form of the bandable structure as follows:
$\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=$ $(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }$
$0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$
Let $h(k)$ be the average of the squares of the $k^{th}$ subdiagonal entries of a covariance:
$h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$
Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$. Figure 1 (a) and (b) show that the $h(k)$ function of the covariance matrix in $\mathcal{U}_{b}$ decays rapidly as the subdiagonal index $k$ increases, while the $h(k)$ function of the covariance matrix in $\mathcal{U}_{bc}$ first decays rapidly, then remains almost unchanged and rises again.
The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart. This may be argued for climate and geospatial applications, where components of the state vectors can be arranged via the geo-distance of the observational sites where the state vectors are associated with. However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class. In such situations, the following thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008b) may be employed. The thresholding covariance class is a class of covariance matrices which are invariant under permutations of components of the state vector $\mathbf{x}_{t}$, defined by
$\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,$ $\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i$
@@
We now present the estimators of the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding based on the sample covariance $\mathbf{S}_{t}^{f}$ of the EnKF ensembles. For any matrix $M=\left[m_{ij}\right]_{p\times p}$, and any $0\leq k<p$, the banding matrix operator with a bandwidth $k$ is
$B_{k}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k)\right].$
The banding estimator of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})$.
The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero. Cai, Zhang and Zhou (2010) updated the banding estimator with a tapering alteration
where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$. It is noted that the tapering estimator is a softer version of the banding estimator by tapering down the level of the thresholding gradually.
In order to provide valid estimation for forecast error covariance belonging to the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, we define a mid-banding operator with bandwidths $k_{1}$ and $k_{2}$,
$B_{k_{1},k_{2}}(M)=\left[m_{ij}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))\right],$
for any matrix $M=\left[m_{ij}\right]_{p\times p}$ and $0\leq k_{1}<(p-k_{2})\leq p$. Then the mid-banding forecast error covariance matrix estimator is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$.
Recall that both the banding and tapering estimators require the variables in $\mathbf{x}_{t}$ to have a natural ordering under a permutation such that the correlation decays as two variables are further apart. For random vectors that cannot be ordered in such a way but belong to the thresholding covariance class $\mathcal{U}_{\tau}\left(q,c_{0}(p),M\right)$, Bickel and Levina (2008b) proposed the thresholding operator
$T_{s}(M)=\left[m_{ij}1\left(\left|m_{ij}\right|\geq s\right)\right],$
at a threshold level $s$. Then the thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})$.
Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$. The selection of these parameters can be carried out by methods in Qiu and Chen (2015). For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in the SM.
To apply the high dimensional estimators of the forecasting error covariance $\boldsymbol{\Sigma}_{t}^{f}$, we need the following distributional assumption.
Assumption 4.
The random vectors $\mathbf{x}_{t,i}^{f}=(x_{t,i1}^{f},\cdots,x_{t,ip}^{f})^{\mathrm{T}}$ are IID and each $x_{t,1j}^{f}\stackrel{{\scriptstyle d}}{{\sim}}F_{t,j}$ and $(x_{t,1j}^{f})^{2}\stackrel{{\scriptstyle d}}{{\sim}}G_{t,j}$, where $G_{t,j}(u)=F_{t,j}(\sqrt{u})-F_{t,j}(-\sqrt{u})$ for $u>0$. There exists a $\lambda_{0}>0$ such that $\max_{1\leq j\leq p}\mathbb{E}(e^{\lambda(x_{t,1j}^{f})^{2}})<\infty\textrm{ for }\lambda\in(-\lambda_{0},\lambda_{0}).$
Assumption 4 effectively assume $(x_{t,1j}^{f})^{2}$ are sub-exponentially distributed, which include the Sub-Gaussian random variables as a special case. The following Proposition 4.1 summarizes the convergence rates of the forecast error covariance matrix estimators by the banding, mid-banding, tapering and thresholding methods.
Proposition 4.1.
Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption 4.
  • (i)
    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$,
    $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (ii)
    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
    $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (iii)
    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$ the optimal rate of convergence of the tapering estimator is
    $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
  • (iv)
    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$ and $n^{-1}{\log p}=o(1)$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$,
    $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
The results in Proposition 4.1 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008a), Cai, Zhang and Zhou (2010) and Bickel and Levina (2008b). And that for the mid-banding estimator in Proposition 4.1 (ii) is a new result whose proof is given in the SM. They imply that the high dimensional covariance estimators are consistent to $\boldsymbol{\Sigma}_{t}^{f}$ under the spectral norm if $\log(p)=o(n)$ or $p=o(e^{cn})$ for any positive constant $c$.
According to Proposition 3.1, the convergence results in Proposition 4.1 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$. Specifically, if $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$ or $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$,
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}\,\text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k}(\mathbf{S}_{t}^{f})\text{ or }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f});\,$ (4.1)
@@
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}\text{ for }\,\hat{\boldsymbol{\Sigma}}_{t}^{f}=\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k);$ (4.2)
[…]
$\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ $\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\ \text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=T_{s}(\mathbf{S}_{t}^{f}).$ (4.3)
These suggest that under the conditions of Proposition 3.1, $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ has the same convergence rate as $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$ for any of the four high dimensional estimators.
@@ 4.2 Convergence of High Dimensional Ensemble Kalman Filter
When the dimension $p$ of the state vector is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 4.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$. Based on the consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman gain matrix $\mathbf{K}_{t}$, we propose the high dimensional EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$. It is noted that for nonlinear model operator, $\mathbf{x}_{t}^{a}$ is defined as (2.6)-(2.7) as indicated before (2.13).
Similar to Theorem 3.1, the following Theorem 4.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
@@
Suppose Assumptions 1(i) or 1(ii), 2, 3 hold and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the conditional distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t,j}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}(\mathbf{x}_{t}^{f},\boldsymbol{\Sigma}_{t}^{f})$ at time $t$. Then, when $p,q\rightarrow\infty$, the forecast and analysis states of the EnKF at time $t$ satisfy
@@
Theorem 4.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (4.1)-(4.3). Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,

### v2

@@ 3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or its poles, the state vectors are highly dependent for the variable components between the two ends of the state vector. Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ function of covariances in $\mathcal{U}_{b}$ is a $L$-type curve as the subdiagonal index $k$ increases, while it is an $U$-type curve for covariances in $\mathcal{U}_{bc}$. For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]$, where $0\leq k_{1}<(p-k_{2})\leq p$.
The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart. However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class. In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}$.
Note that all of the banding, mid-banding, tapering and thresholding estimators depend on either the bandwidth $k$ or the threshold $s$ which can be selected by methods in Qiu and Chen (2015). For the bandwidths $k_{1},k_{2}$ of the mid-banding estimator, we propose a similar selection method and details can be found in Section S.2 of the SM.
The following Proposition 3.2 summarizes the convergence rates of the forecast error covariance estimators by the banding, mid-banding, tapering and thresholding methods.
Proposition 3.2.
Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption S.3.1 of the SM.
  • (i)
@@
    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (ii)
    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
  • (iii)
    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
  • (iv)
    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
The results in Proposition 3.2 (i), (iii) and (iv) are, respectively, from Bickel and Levina (2008b), Cai et al. (2010) and Bickel and Levina (2008a). And that for the mid-banding estimator in Proposition 3.2 (ii) is a new result whose proof is given in the SM. They imply that the high dimensional covariance estimators are consistent to $\boldsymbol{\Sigma}_{t}^{f}$ under the spectral norm if $\log(p)=o(n)$ or $p=o(e^{cn})$ for any positive constant $c$.
According to Proposition 3.1, the convergence results in Proposition 3.2 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$. Specifically,
@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
When the dimension $p$ of $\mathbf{x}_{t}$ is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 3.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$. Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$. To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
@@
Suppose Assumptions 1(i) or 1(ii), 2, 3 hold and the forecast ensemble $\{\mathbf{x}_{t,j}^{f}\}_{j=1}^{n}$ are independently sampled from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t}^{f}$ and covariance $\boldsymbol{\Sigma}_{t}^{f}$. Then, the forecast and analysis states of the EnKF at time $t$ satisfy
@@
Theorem 3.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (3.1). Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
@@
For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ as $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM. Furthermore, Proposition S.3.1 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic distributions of $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.